OpenStudy (anonymous):

Write the exponential expression using radicals. 19^-7/2

OpenStudy (anonymous):

\[19^{-7/2} = (19^{-7})1/2 = \sqrt{19^{-7}} = \sqrt{1/19^{7}}\]